Chapter 1

A sample document

1.1 1st language: English

This document illustrates some points of the package polyglot. For instance, the British date is 8th June 1999.

The American one is June 8, 1999.

\uselanguage[date]{american}{\today}

1.2 2^{e} langue : français

The first things you will notice in French is the indentation of the first line after the title, which is in the layout group, and the small space just before the colon. Other features are :

– Dashes in every level of itemize environment.

– Ligatures for a lot of accents (and encyclopædic task). encyclop"aedic

– French guillemets . << guillemets >>.

– French spacing.

Today is 8 juin 1999.

1.3 3.^{er} idioma: español

Also ligatures, guillemets, new layout, etc. Today is 8 de junio de 1999.

1.4 4. Sprache: Deutsche

The usual ligatures are implemented, as German quotes“ "`German quotes"', "ck, usw. In OT1 you get low umlauts: ä, ö, ü. Today is 8. Juni 1999.

1.5 5th language: hebrew

In this section we illustrate the use of the package for romanizing languages. In r_hebrew we define some ligatures which simplifly the typing of text (following the ISO recomendations, with lots of diacritic marks). For example:

ba.r{ā \atop .}, timz̄¸ēn{ä \atop .}, tez̄ē, rōš, ya.vô', ka.lû', na.vî', pa.r{ä \atop .}, z¸ë, qĕnë, ga.lö,
ba.n\begin{matrix}a \\ . \end{matrix}(y)w, rĕ'iyy{ä \atop .}, ha.-a.rȩz̄, wȩ-¸ĕd̄ôm, rûah.

b.ar-.a, tim-z-.en".a, te-z-e, r-o^s, y.av^o'',
k.al^u'', n.av^i'', p.ar".a, z".e, q'en"e,
g.al"o, b.an.a(y)w, r'e''iyy".a, h.a--.ar.e-z,
w.e--'.e-d^om, r^ua.h